OpenStudy (anonymous):

What is the equation of a line that passes through the point (-3, 1) and is parallel to a line with a slope of 2? y = 2x + 1 y = 2x + 4 y = 2x + 5 y = 2x + 7

OpenStudy (traceur55):

The equation would probably be found someone a little more suited to do so. Perhaps the Mathematics section?

OpenStudy (traceur55):

But, for real. Let's break this down. Y=Mx+b Y is 1. M is 2(Slope) X is? You guessed it -3. So your equation should look like. 1=2(-3)+b Now: 1=-6+b. You can finish.

OpenStudy (anonymous):

k